OpenStudy (anonymous):

Help check my work please: Box A has a volume of 32 cubic meters. Box B is similar to Box A. To create Box B, Box A's dimensions were tripled. What is the volume of Box B? 864 m3 288 m3 96 m3 32 m3 I got 864 the first time, but then the second time I got 96. Help please!

OpenStudy (anonymous):

@ganeshie8 please help me

hartnn (hartnn):

when dimensions are multiplied by 'x' Area gets multiplied by \(\large x^2\) and volume gets multiplied by \(\huge x^3 \)

OpenStudy (anonymous):

Okay, I'm still confused. The first time I did 3^3 = 27. 32 * 27 = 864 but then when I checked it again I did it like this (1/3)^3 = 32/ V 1/27 = 32/V V=96 and i got a different answer. Both are choices

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

32*27 = 864m^3 is correct answer.

hartnn (hartnn):

how did you check it ?

OpenStudy (anonymous):

I just redid the problem again to make sure, are you sure that 864 is correct? I'm so confused

OpenStudy (anonymous):

You were correct, thank you so much for helping me

hartnn (hartnn):

1/27 = 32/V gives V = 27*32 =864 and welcome ^_^